Calculate the number of grams of copper produced in 5.67 hours by the electrolysis of molten CuCl2 if the electrical current is 11.7 A.

You may assume the molar mass of elemental copper is 63.546 g/mol.
Answer #1

Electrolysis equation is:

Cu2+ + 2e- ------> Cu

1 mol of Cu requires 2 mol of electron

1 mol of electron = 96485 C

So,1 mol of Cu requires 192970 C

let us calculate the charge passed:

t = 5.67 hr = 5.67*3600 s = 2.041*10^4 s

time, t = 2.041*10^4 s

Q = I*t

= 11.7 A * 2.041*10^4 s

= 2.388*10^5 C

mol of Cu plated = 2.388*10^5/192970 = 1.238 mol

Molar mass of Cu = 63.546 g/mol

mass of Cu = number of mol * molar mass

= 1.238 * 63.546

= 78.6 g

Answer: 78.6 g